The perimeter of a rhombus is 40cm, and one diagonal is 12cm long. How long is the other diagonal?
Answer by venugopalramana(3286)

The perimeter of a rhombus is 40cm, and one diagonal is 12cm long. How long is the other diagonal?
RHOMBUS PERIMETER =4*SIDE=40
SIDE=40/4=10
DIAGONALS BISECT EACH OTHER AT RIGHT ANGLES IN RHOMBUS.HENCE
SIDE^2 = (D1/2)^2+(D2/2)^2
10*10 = (12/2)^2+(D2/2)^2
(D2/2)^2=100-36=64
D2/2=8
D2=16